A 1.42 g sample of a pure compound, with formula M
Answer #1

BaSO4 precipitate formed

means all sulfate combines with Ba

233.43 g BaSO4 contain 96.06 g SO4-2

2.33 g BaSO4 contain 2.33 x 96.06 / 233.43 = 0.96 g SO4-2

means our M2SO4 contain 0.96 g SO4-2

mass of M2 = 1.42 - 0.96 = 0.46 g

0.96 g SO4-2  combines with 0.46 g M2

96.06 g SO4-2  combines with 96.06 x 0.46 / 0.96 = 46.03 g M2

so mass of M = 46.03 / 2 =23 g

Na atomic weight = 23 g /mol

so M = Na